OpenStudy (anonymous):

What is the x-coordinate for the minimum point in the function f(x) = 4 cos(2x − π) from x = 0 to x = 2π? x = pi over 2 x = pi over 4 x = 2π x = π

OpenStudy (anonymous):

@aloud @phi @mathmate

OpenStudy (phi):

do you know what cos looks like from 0 to 2pi? https://www.google.com/search?q=cos%28x%29&gbv=2&sei=FSSEVeKzLIevsAWqqoKYAQ

OpenStudy (anonymous):

i thought it intercepts at -4

OpenStudy (phi):

the smallest value will be y= -4 for a "normal cosine" from 0 to 2pi, the minimum value is at pi (exactly half way from 0 to 2pi)

OpenStudy (anonymous):

no I dont know how to graph them that well

OpenStudy (phi):

so you want to find where the "angle" of your problem f(x) = 4 cos(2x − π) from x = 0 to x = 2π equals pi in other words, what is "x" so that 2x - pi = pi

OpenStudy (anonymous):

would you plug in the x=0 for the x in 2x-pi=pi

OpenStudy (anonymous):

like substitution

OpenStudy (phi):

but using algebra 2x - pi = pi add +pi to both sides 2x - pi + pi = pi+pi on the left -pi+pi add up to zero. on the right you get 2pi 2x = 2pi now divide both sides by 2

OpenStudy (anonymous):

it would be x=pi because the 2's would cancel right?